# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v5

@@ 4.1. AI-Assisted Flight Booking
After collecting data from 500 participants, the researchers model the percentage of trials where the participants incorrectly follow the AI’s advice. They find a statistically significant difference between the AI explanation condition and the control condition. The researchers interpret this as evidence that AI explanations promote over-reliance.
@@
How would a rational participant form beliefs about the posterior probability of the state using Equation 1? Participants are not provided with information about the prior probability of the state. This would be acceptable if the signal provided them with posterior beliefs about the uncertain state. However, the signal does not provide the required posterior probability: recall that participants receive the estimated probability that the model is correct conditional on the features, , but not conditional on the model’s prediction. They therefore cannot obtain sufficient information from the data-generating model to guide their choice of action.
@@ 4.2. Election Forecast Displays
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to the closeness of the race, defined as the difference between the predicted vote share of the first candidate and 0.5. The researchers postulate that a better display induces more sensitivity. Using statistical significance testing, they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types. Using posthoc comparisons, they declare that the animated graphic leads to better voting decisions because it induces the most sensitivity.
@@
5. Applying the Normative Framework to HCAI Studies
To illustrate the applicability of our definition, we present a meta-study in which we analyzed a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions without a well-defined decision problem.
@@ 5. Applying the Normative Framework to HCAI Studies
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable: Does the task given to participants admit a ground truth state and induce uncertainty about the state, at least across participants? Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way? We identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information for a rational agent to best respond to the task. Finally, we noted whether the researchers reported conclusions using a different scoring rule than the one provided to participants, and whether participants were incentivized using a flat scoring rule (i.e., a rule that does not distinguish between responses based on their distance to the ground truth). Note that studies that pay participants a flat reward require that researchers report conclusions using a different scoring rule.
In some cases, coding was challenging because the paper did not provide much detail on what exactly was communicated to participants, nor provide the study interface as a supplement. We noted such cases.
@@
Results. Of the 46 studies examined, 7 (15%) studied tasks to which the framework could not be applied. These included studies relegated to perceptions (e.g., of fairness, accuracy) and of subjective appraisals of music or movie recommendations or emotion. One study posed questions and drew conclusions about the diagnostic utility of AI recommendations, but elicited only perceived utility (Cai et al., 2019). We coded this example as Applicable.
@@
All 39 of the remaining studies drew conclusions about flaws in human decisions, such as by pointing to over- or underreliance on the AI or speculating about why participants failed to get the correct answer. However, participants were only given sufficient information to identify the best decision response across all conditions in 7 of these 39 (18%), and across some conditions in another 3 of 39 (7.7%). In the latter cases, informational differences between conditions that the researchers compared could lead to different a priori expectations of decision quality, yet these differences were not accounted for in comparing participants’ behavior between conditions555The calculation of separate performance benchmarks would address such ambiguity (Wu et al., 2023).. There was not enough information to determine if participants were presented with a well defined problem in 4 (10%) of the remaining studies.
@@
Only 7 of the 39 (18%) studies to which the framework is applicable used the same scoring rule to analyze the responses as was used to incentivize participants. 25 (64%) did not, one did in at least one experiment, and 6 others (15%) did not report enough detail for us to know. Finally, 14 (36%) of the 39 used a flat rule for incentives, one did in some experiments, 16 (36%) did not, and 7 (18%) did not report enough detail for us to know.
@@
We observed that even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent. As described above, this misinterprets the performance of participants guided by a different scoring rule. In some cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model. The most common scenario was when the researchers gave participants information about the model’s held-out accuracy but the participants experienced instances where the model had a different accuracy. When the study design includes feedback, this leads to a contradiction between the instructions and what participants observe which makes it difficult to draw evaluative claims about behavior.
6. Recommendations for Practice
The lack of consensus around how to define a decision problem and the minimum required components for normative analysis leads to research landscape in which researchers can easily be tempted to overinterpret the results of their studies. Readers may not recognize the overinterpretations, leading to a research literature based on invalid claims about human biases. Adopting a more rigorous standard in the form of the proposed framework, based in established theoretical foundations for decision making from informative signals (Savage, 1972; Bergemann and Morris, 2019), stands to improve empirical decision research in several ways. First, it improves the validity of interpretations of individual studies. The variation induced when a decision problem is not clearly reported nor communicated to participants is consequential to interpretation in that it cannot be separated from variation induced by true differences in how people do a task under different conditions. Identifying the key sources of loss—much less separating them given experiment results—is a hard problem made impossible. Adopting a standard where the normative decision arises from a well-defined process stands to improve researchers’ ability to point to possible sources of performance loss, and raise their awareness of the limitations on what be concluded from study results when a decision problem is under-communicated. Second, it paves the way for more rigorous theory development. Only when researchers define the task they study as a complete decision problem can the results of experiments that study slightly different problems (e.g., use different data-generating models) be meaningfully contrasted to drive the development of theory.
Consequently, researchers wishing to design a normative decision experiment should communicate all necessary components of the decision problem both to study participants and in publications. Specifically, we recommend that researchers:
Clearly convey action and state spaces and a scoring rule. The finite action space and statespace should be clearly communicated to participants, and the action space reflected in the instruments used to collect participant responses. Because the optimal action can only be determined relative to a scoring rule, participants should be given a scoring rule to reduce heterogeneity in their choice of action, even if the experimenter feels comfortable assuming that the participants will try their best and does not use the scores in assigning monetary rewards.
Provide prior and/or sufficient information about the data-generating model to calculate posterior. Explicit specification of priors is rare in decision studies used to evaluate interfaces in practice. As discussed above, sufficient information about the data-generating model must be provided to arrive at the Bayesian posterior distribution. Whenever the signal on its own does not induce a posterior, the prior should be endowed with sufficient information about should be provided to enable estimating and . Exemplar-based training procedures such as have been used to teach humans when to defer to an AI model can be employed for this purpose . More generally, using presentation formats that are more likely to be automatically processed (e.g., conveying distribution information via frequency animations (Goldstein and Rothschild, 2014; Hullman et al., 2015)) or validating participants’ priors through elicitation (e.g., (Kim et al., 2019, 2020)) can help researchers understand and minimize prior loss.
Explicitly motivate scoring rules for incentivization and evaluation. Beyond defining the components of the target decision problem, a next step is for researchers to motivate the problem definition they use. Recent work describes the use of a rational agent framework similar to the one presented here to analyze whether a decision study is likely to provide agents with a sufficient incentive to motivate them to consult the signals (Wu et al., 2023). Theory around scoring rules (Li et al., 2022) and what statistical quantities are elicitable (Lambert, 2011) stand to offer prescriptive advice if incorporated in human decision research.
Compare to best attainable performance. We observed that even when researchers invest effort to design scoring rules to incentivize participants, they do not often compare the observed behavioral performance to the best attainable performance under the rule. Researchers should default to analyzing results in the space of the scoring rule used for incentives, and provide a rationale whenever they deviate. How to estimate sources of loss in an unbiased way is a rich area for future work.
Elicit measures to estimate specific sources of loss. An important area for future work is to establish under what conditions the losses we conceptualize can be estimated. Separating losses entails adding carefully constructed questions to the experimental procedure, and/or constructing benchmarks based in rational agent performance to separate sources of loss in observed results. For example, Wu et al. (2023) use the notion of a calibrated behavioral agent along with a rational agent benchmark and baseline (expected performance of a rational agent under only the prior) to separate optimization and receiver loss.
Conduct design analysis. Researchers doing normative decision research can benefit from conducting analyses prior to running the study which simulate the idealized actions of an agent to vet the experiment design (see, e.g., Wu et al. (2023)’s pre-study analyses for visualization studies and Guo et al. (2024)’s analysis of the potential for complementarity in AI-assisted decisions). Simulation-based design analysis forces the researcher to explicitly define the parameters of their decision problem, which can elucidate what information needs to be communicated to a participant in order to enable an optimal response. It also enables the researcher to test variations on their data-generating model and scoring rule, which can help the researcher realize under what conditions (1) the experiment can have acceptable statistical power and (2) participants can be paid fairly. Design analysis helps the researcher choose conditions (e.g., AI explanations) that provide informative signals for the decision task, narrowing down what kinds of research questions can be answered using a particular decision problem. For example, at a minimum, to be informative of how effectively people use signals, a study design must have sufficient statistical power to separate the rational agent performance with and without the signal.
6.1. Discussion
By arguing that researchers adopt and sufficiently communicate to participants a widely-applicable definition of a decision problem, our goal is to improve the integrity of claims made from human subjects experiments. We now address how the proposed normative decision framework sheds light on fundamental uncertainty in generalizing study results to real-world contexts. We discuss several common misunderstandings of a rational agent approach to defining normative decisions. We conclude with limitations and directions for future work.
6.2. Goals and Incentives in the Lab versus the World
@@
When participants are not given ample information to “solve” at least in principle the decision problem, we cannot bring well-formed expectations to the results, and should expect heterogeneity in the behaviors they exhibit. Such “heterogeneity by design” renders the results uninterpretable for evaluative purposes like identifying forms of bias.
@@
This requirement of communicating to study participants sufficient information for a rational agent to optimize may seem unnecessarily strict if one assumes that participants’ domain knowledge alone should enable them to evaluate the information about the state captured in signals. For example, in the flight booking example, a researcher may be tempted to assume that even without being given any information about the probability that the model prediction is correct, a participant will be able to use their prior knowledge about features of delayed flights to infer the probability that the model prediction is correct. The problem is that once a researcher makes this assumption, they are no longer evaluating participants on only the information they provided. They are also implicitly evaluating them on how well their prior beliefs about the state and the validity of the model’s learned information align with the researcher’s choice of data-generating model. There are various implications of this misalignment that we may find problematic upon reflection. For example, under such conditions, participants must bring identical beliefs to the researchers’ in order to have a chance of scoring well.
@@
6.3. Common Misunderstandings
Our argument is that the interpretation of results is restricted when a well-defined decision problem is not conveyed to participants, not that researchers cannot conduct comparisons at all without a well-defined decisions. Researchers can conduct and report on studies that do not conform to this approach, but they cannot conclude that people made bad decisions if they were not given sufficient information to know what a good decision is. For example, imagine a study comparing the decision performance of users with and without the use of an AI, where no prior is given to participants, and no scoring rule is given. Imagine that performance in the study is measured in terms of accuracy, and average accuracy is higher among participants who had access to the AI. Given an appropriately designed randomized experiment, the researcher could conclude that there is a causal effect of the AI on decisions. However, they could not claim that people are making worse or biased decisions without the AI, because they are scoring their decisions using information they never provided to participants. When we say that one condition did better, we should acknowledge that we did not give them a chance to fully understand and optimize their behavior for the problem that we analyzed.
Finally, a common misinterpretation of the use of a rational agent model is that we are implying that human behavior is expected to be perfectly rational, or approximately rational. This is not the case. While a rational agent approach formally justifies what it means to best use information to inform the selection of an action using statistical decision theory, participants in our experiment may use different strategies. The point is that without a definition of what it means to best respond to the information we have provided, we are hard-pressed to evaluate participant behavior. Even if we do not expect this procedure to characterize what human agents do, it provides a basis for interpreting the results in terms of deviations from a well-understood standard.
@@ 6.4. Limitations & Future Work
Some have voiced pessimistic views on the value of monetary incentives for experiment participants (e.g., (Read, 2005)). As Wu et al. (2023) show, this view is to be expected under a rational agent approach whenever the value of information to the decision problem (i.e., the difference in the rational agent’s expected score given only the prior versus with access to the signals (Wu et al., 2023)) is small. More broadly, the insensitivity of human participants to monetary incentives are orthogonal to the crux of our argument, which is philosophical and concerned with what is knowable more so than the pragmatics of how much behavior changes when a problem is well-defined.
Scoring rules can sometimes be challenging for study participants to interpret; theoretical incentive compatibility does not guarantee understanding on the part of subjects. There is a rich literature discussing possible descriptions that can be used with common rules like the quadratic scoring rule (e.g., (Artinger et al., 2010)). Interfaces researchers may be in a good position to extend this literature through careful observational studies.
Ultimately, the goal of our work is to give researchers the tools to recognize that when we do experiments, we are restricted to learning about behavior in the artificial worlds we create. As much as we might want to equate our results with some real world setting used as motivation, extrapolating from the world of the controlled experiment to the real world will always be a leap of faith. It is therefore the responsibility of the experimenter to fully understand their experimental world, and, for the purposes of offering the results as a form of potentially generalizable knowledge, to give their participants a chance to do so as well.

### v6

@@ 4.1. AI-Assisted Flight Booking
After collecting data from 500 participants, the researchers model the percentage of trials where the participants incorrectly follow the AI’s advice. They find a statistically significant difference between the AI explanation condition and the control condition, with the rate of incorrectly following the AI’s advice being higher in the former. The researchers interpret this as evidence that AI explanations promote over-reliance.
@@
How would a rational participant form beliefs about the posterior probability of the state? Participants are not provided with information about the prior probability of the state. This would be acceptable if the signal provided them with posterior beliefs about the uncertain state. However, the signal does not provide the required posterior probability: recall that participants receive the estimated probability that the model is correct conditional on the features, , but not conditional on the model’s prediction. They therefore cannot obtain sufficient information from the data-generating model to guide their choice of action.
@@ 4.2. Election Forecast Displays
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to a variable denoting the closeness of the race, operationalized as the difference between the predicted vote share of the first candidate and 0.5. The researchers postulate that a better display induces more sensitivity. Using statistical significance testing, they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types. Using posthoc comparisons, they declare that the animated graphic leads to better voting decisions because it induces the most sensitivity.
@@
5. Applicability of the Normative Framework to HCAI Studies
To illustrate the applicability of our definition, we present a meta-study in which we analyzed a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions without a well-defined decision problem. Note that HCAI studies are just one example of the kinds of decision studies to which our work applies (Section 2.3).
@@ 5. Applicability of the Normative Framework to HCAI Studies
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable: Does the task given to participants admit a ground truth state and induce uncertainty about the state, at least across participants? Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way? We identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information for a rational agent to best respond to the task. Finally, we noted whether the researchers reported conclusions using a different scoring rule than the one provided to participants, and whether participants were incentivized using a flat scoring rule (i.e., a rule that does not distinguish between responses based on their distance to the ground truth). Note that studies that pay participants a flat reward require that researchers report conclusions using a different scoring rule, introducing confounding.
In some cases, coding was challenging because the paper did not provide much detail on what exactly was communicated to participants, nor provide the study interface as a supplement. This demonstrates a lack of sensitivity on the researchers’ part when it comes to providing requisite information to interpret their results. We noted such cases, and discuss how specific omissions are problematic.
@@
Results. Of the 46 studies examined, 39 (85%) studied tasks to which the framework was applicable. One study posed questions and drew conclusions about the diagnostic utility of AI recommendations, but elicited only perceived utility (Cai et al., 2019). We coded this example as Applicable. The remaining 7 (15%) included studies relegated to perceptions (e.g., of fairness, accuracy) and of subjective appraisals of music or movie recommendations or emotion.
@@
All 39 of the applicable studies drew conclusions about flaws in human decisions, such as by pointing to over- or underreliance on the AI or speculating about why participants failed to get the correct answer. However, when we evaluated whether participants were given sufficient information to identify the best decision response across all conditions, this was true in only 7 of the 39 (18%) cases. In another 3 of 39 (7.7%), participants were given sufficient information to identify the best decision response in some subset of conditions. In the latter cases, informational differences between conditions that the researchers compared could lead to different a priori expectations of decision quality, yet these differences were not accounted for in comparing participants’ behavior between conditions555The calculation of separate performance benchmarks would address such ambiguity (Wu et al., 2023).. While the framework was clearly applicable, there was not enough information to determine if participants were presented with a well defined problem in 4 (10%) of the remaining studies. In other words, at best, over 70% of the surveyed studies did not give participants an opportunity to do well by the evaluation approach the researchers applied.
@@
Only 7 (18%) of the 39 studies to which the framework is applicable used the same scoring rule to analyze the responses as was used to incentivize participants, illustrating a lack of awareness of the implications of using different rules for interpreting the meaning of responses. 25 (64%) did not, one did in at least one experiment, and another 6 (15%) did not report enough detail for us to know. Finally, 14 (36%) of the 39 used a flat rule for incentives, one did in some experiments, 16 (36%) did not, and 7 (18%) did not report enough detail for us to know.
@@
We observed that even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent. In some cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model, such that the results interpretation did not hold. The most common scenario was when the researchers gave participants information about the model’s held-out accuracy but the participants experienced instances where the model had a different accuracy. Especially when the study design includes feedback, this leads to a contradiction between the instructions and what participants observe which makes it difficult to draw any evaluative claims about behavior.
Altogether, these results suggest that the framework is widely applicable to swaths of studies currently being conducted in HCI and related fields. Moreover, they suggest that researchers stand to improve the interpretability of their research by adopting it by following recommendations that it implies, as we describe further below.
6. Discussion: Goals and Incentives in the Lab versus the World
@@
By arguing that researchers adopt and sufficiently communicate to participants a widely-applicable definition of a decision problem, our goal is to improve the integrity of claims made from human subjects experiments. When participants are not given ample information to “solve” at least in principle the decision problem, we cannot bring well-formed expectations to the results, and should expect heterogeneity in the behaviors they exhibit. Such “heterogeneity by design” renders the results uninterpretable for evaluative purposes like identifying forms of bias.
@@
6.1. Implications for Research Ethics
The requirement of communicating to study participants sufficient information for a rational agent to optimize may seem unnecessarily strict if one assumes that participants’ domain knowledge alone should enable them to evaluate the information about the state captured in signals. For example, in the flight booking example, a researcher may be tempted to assume that even without being given any information about the probability that the model prediction is correct, a participant will be able to use their prior knowledge about features of delayed flights to infer the probability that the model prediction is correct. The problem is that once a researcher makes this assumption, they are no longer evaluating participants on only the information they provided. They are also implicitly evaluating them on how well their prior beliefs about the state and the validity of the model’s learned information align with the researcher’s choice of data-generating model. There are various implications of this misalignment that we may find problematic upon reflection. For example, under such conditions, participants must bring identical beliefs to the researchers’ in order to have a chance of scoring well. A related point has been made in research on crowdsourcing, where it has been observed that there are often many plausible ways to interpret a crowdsourced task based on the information workers are provided, but work requesters often use quality control algorithms that recognize and pay only one of these (alkhatib2019street; kairam2016parting).
@@ 6.2. Limitations & Future Work
Our argument is that the interpretation of results is restricted when a well-defined decision problem is not conveyed to participants, not that researchers cannot conduct comparisons at all without a well-defined decisions. Researchers can conduct and report on studies that do not conform to this approach, but they cannot conclude that people made bad decisions if they were not given sufficient information to know what a good decision is. For example, imagine a study comparing the decision performance of users with and without the use of recommendations from an AI, where no prior is given to participants, and no scoring rule is given. Imagine that performance in the study is measured in terms of accuracy, and average accuracy is higher among participants who had access to the AI. Given an appropriately designed randomized experiment, the researcher could conclude that there is a causal effect of the AI on decisions. However, they could not claim that people are making worse or biased decisions without the AI, because they are scoring their decisions using information they never provided to participants. When we say that one condition did better, we should acknowledge that we did not give them a chance to fully understand and optimize their behavior for the problem that we analyzed.
A common misinterpretation of the use of a rational agent model is that we are implying that human behavior is expected to be perfectly rational, or approximately rational. This is not the case. While a rational agent approach formally justifies what it means to best use information to inform the selection of an action using statistical decision theory, participants in our experiment may use different strategies. The point is that without a definition of what it means to best respond to the information we have provided, we are hard-pressed to evaluate participant behavior. Even if we do not expect this procedure to characterize what human agents do, it provides a basis for interpreting the results in terms of deviations from a well-understood standard.
@@
Some have voiced pessimistic views on the value of monetary incentives for experiment participants (e.g., (Read, 2005)). As Wu et al. (2023) show, this view is to be expected under a rational agent approach whenever the value of information to the decision problem (i.e., the difference in the rational agent’s expected score given only the prior versus with access to the signals (Wu et al., 2023)) is small. This motivates analyzing whether a decision study is likely to provide agents with a sufficient incentive to motivate them to consult the signals (Wu et al., 2023).
Of course, scoring rules can sometimes be challenging for study participants to interpret. Theoretic incentive compatibility does not guarantee understanding on the part of subjects. There is a rich literature discussing possible descriptions that can be used with common rules like the quadratic scoring rule (e.g., (Artinger et al., 2010)). Interfaces researchers may be in a good position to extend this literature through careful observational studies. More broadly, however, the insensitivity of human participants to monetary incentives are orthogonal to the crux of our argument, which is philosophical and concerned with what is knowable more so than the pragmatics of how much behavior changes when a problem is well-defined.
7. Conclusion
We present definitions of a decision problem and normative response synthesized from statistical decision theory and information economics for the purpose of improving the interpretability of decision studies in HCI, visualization, human-centered AI, and related fields. We demonstrate the applicability of the framework to recent HCAI studies, and provide specific recommendations to researchers and examples of how confounded study designs can be made interpretable or well-formed within the framework. Ultimately, the goal of our work is to give researchers the tools to recognize that when we do experiments, we are restricted to learning about behavior in the artificial worlds we create. As much as we might want to equate our results with some real world setting used as motivation, extrapolating from the world of the controlled experiment to the real world will always be a leap of faith. It is therefore the responsibility of the experimenter to fully understand their experimental world, and, for the purposes of offering the results as a form of potentially generalizable knowledge, to give their participants a chance to do so as well.